Numerical Analysis I, Fall 2010 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 02

Assigned Sep 25, 2010.

1. Section 1.3: Problems 3, 4, 7, 14, 15.

2. Section 2.1: Problems 12(*), 18.